OpenStudy (anonymous):

I need help solving a problem out of my book: The SI unit of length is the meter, which is defined as the length equal to 1650763.73 wavelengths of the light emitted by a particular energy transition in krypton atoms. Calculate the frequency of the light to three significant figures.

OpenStudy (anonymous):

Got it: 1650763.73 wavelengths per second times 3.0x10^8 m/s (speed of light) = 4.95x10^14 which is the frequency.

OpenStudy (aaronq):

Thats right, i did it the long way to double check it. \(\lambda =\dfrac{1~m}{1650763.73~oscillations}=6.06*10^{-7} m\) \(\nu*\lambda=c \rightarrow \nu=\dfrac{c}{\lambda}=\dfrac{3.0*10^8~m/s }{6.06*10^{-7} m}=4.95*10^{14}~Hz\)

OpenStudy (vincent-lyon.fr):

By the way, this is not the present definition of the metre, but the previous one. Today 1 metre is the distance travelled by light in a vacuum in 1/299,792,458 s.